Iodine-131 is a radioisotope with a half-life of 8 days and can be used in the treatment of thyroid cancer. Calculate the percentage of iodine-131 that would have decayed after 24 days


1
Expert's answer
2021-01-21T03:49:29-0500

"N=N_0(\frac{1}{2})^{\frac{t}{t_{1/2}}}" , where N is the amount at given time (in this case - the percentage remaining), N0 is the initial amount (in this case it is equal to 100 percent), t is time, and t1/2 is the half-life. Therefore,


"N=100\%\times(\frac{1}{2})^{\frac{24days}{8days}}=12.5\%" (remaining).

Hence 100 - 12.5 = 87.5% would have decayed.


Answer: 87.5%